### v1

# Do Pre-Trained Language Models Detect and Understand Semantic Underspecification? Ask the DUST!

Abstract
In everyday language use, speakers frequently utter and interpret sentences that are semantically underspecified, namely, whose content is insufficient to fully convey their message or interpret them univocally. For example, to interpret the underspecified sentence “Don’t spend too much”, which leaves implicit what (not) to spend, additional linguistic context or outside knowledge is needed. In this work, we propose a novel Dataset of semantically Underspecified Sentences grouped by Type (DUST) and use it to study whether pre-trained language models (LMs) correctly identify and interpret underspecified sentences. We find that newer LMs are reasonably able to identify underspecified sentences when explicitly prompted. However, interpreting them correctly is much harder for any LMs. Our experiments show that when interpreting underspecified sentences, LMs exhibit little uncertainty, contrary to what theoretical accounts of underspecification would predict. Overall, our study reveals limitations in current models’ processing of sentence semantics and highlights the importance of using naturalistic data and communicative scenarios when evaluating LMs’ language capabilities.
Do Pre-Trained Language Models Detect and Understand Semantic Underspecification? Ask the DUST!
Frank Wildenburg College of Informatics University of Amsterdam f.c.l.wildenburg@uva.nl                        Michael Hanna ILLC University of Amsterdam m.w.hanna@uva.nl                        Sandro Pezzelle ILLC University of Amsterdam s.pezzelle@uva.nl
1 Introduction
Speakers can almost effortlessly deal with semantic underspecification, a widespread phenomenon that occurs when a linguistic signal does not fully convey all the information required for communication to succeed Frisson (2009); Harris (2020). This is possible because, in a normal state of affairs, humans have access to further linguistic or extra-linguistic information coming from the conversation, the surrounding context, or shared knowledge. If this is the case, speakers will have no trouble understanding the sentences “I’ll meet you there” or “I saw you on the hill with the telescope”, even though these sentences leave underspecified where “there” is, or which interlocutor had “a telescope”. It has been argued that underspecification and the related phenomenon of ambiguity are not a hindrance in language, but rather a desirable feature of human language communication Piantadosi et al. (2012); they allow for more concise utterances, which makes language more efficient. Indeed, humans excel at making inferences (Grice, 1969), which is less cognitively taxing compared to articulating speech Levinson (2000).
S1 Don’t spend too much.
S2 Don’t spend too much cash.
Exp. 1 Is S2 more specified than S1?
S1 The bag is on the chair. It is green
S2 The bag is on the chair. The chair
is green.
Exp. 2 Does S1 mean the chair is green? ?
Does S2 mean the chair is green?
Does S1 mean the chair isn’t green? ?
Does S2 mean the chair isn’t green?
Figure 1: In experiment 1 (top), we test if LMs distinguish underspecified sentences (S1) from more specified counterparts (S2) when explicitly prompted. In experiment 2 (bottom), we test if LMs interpret correctly S1 and S2 in a more naturalistic communicative setting.
While humans are good at dealing with semantically underspecified language by leveraging additional information, how pre-trained transformer language models (LMs) behave when faced with this phenomenon is an open question. Despite the growing literature exploring the semantic capabilities of last-generation LMs Ettinger (2020); Rogers et al. (2020); Hanna et al. (2023), little attention has been paid to this problem. Furthermore, the few previous works generally did not distinguish between the various facets of ambiguity and underspecification Liu et al. (2023) or only focused on ambiguity Stengel-Eskin et al. (2024); Ortega-Martín et al. (2023); Fantechi et al. (2023).
However, handling semantically underspecified language is critical for LMs. Since underspecified sentences can license multiple interpretations, choosing one arbitrarily can lead to undesired or even harmful consequences for communication (see Hutchinson et al., 2022; Pezzelle, 2023, for an in-depth discussion of this issue). Thus, LMs should (1) recognize semantically underspecified inputs and (2) interpret them appropriately, ideally with no biases toward a default reading.
We build on and extend previous work by asking two new questions: (1) to what extent can LMs detect if a sentence is (under)specified? (2) How do LMs interpret such underspecified sentences compared to more specified counterparts? To this end, we introduce DUST [^1] We release DUST and code for our experiments here: https://github.com/frank-wildenburg/DUST, a Dataset of semantically Underspecified Sentences (and more specified counterparts) grouped by the Type of underspecification they belong to, and propose a suite of experiments to answer the questions above using DUST. To categorize instances of underspecification, we build on Egg’s (2010) proposed taxonomy.
Our experiments and analysis show that (1) distinguishing semantically underspecified sentences from specified counterparts is not a trivial task for current LMs. While newer, better-performing models achieve reasonable performance when given explicit instructions, other models fare only slightly better than random. Moreover, (2) when asked to interpret underspecified sentences in a more naturalistic communicative scenario without explicit guidance, all LMs fall into the trap of interpreting them similarly to their more specified versions. This suggests that, in the absence of a specific prompt, these models assign biased or default interpretations to underspecified and ambiguous sentences.
Our findings confirm that current LMs, including the best-performing Llama 2 Touvron et al. (2023) and Mistral Jiang et al. (2023), struggle with underspecified and ambiguous language (in line with Liu et al., 2023). This reveals more general limitations with the processing of sentence semantics. Moreover, our study highlights the importance of methodological choices, such as experimental setting, or the level of informativeness of prompts, in fine-grained evaluations of LMs’ capabilities.
2 Related work
2.1 Semantic Underspecification
Semantic underspecification is a phenomenon in which the semantic material of a sentence “leaves open” possibilities for readers of the text that may then be “filled in” through non-linguistic information Zwicky and Sadock (1975); Frisson (2009); Egg (2010); Harris (2020). The phenomenon has traditionally been studied through the lens of (formal) linguistics and semantics (Zwicky and Sadock, 1975; Lappin, 2000; Egg, 2010; Harris, 2020, inter alia), although it has also been studied in other fields, such as the neuroscience of language processing Frisson (2009) and information theory Piantadosi et al. (2012); Franzon and Zanini (2022).
Semantic underspecification is often studied in tandem or in contrast with ambiguity, a related phenomenon. The difference between the two is that underspecified sentences sometimes have only one reading (which may be clarified by non-linguistic information), whereas ambiguous sentences have multiple readings Zwicky and Sadock (1975). However, all ambiguous sentences are also underspecified; after all, some non-linguistic information will disambiguate between readings. Hence, underspecification can be seen as a generalization of ambiguity. Despite this, the terms are sometimes used interchangeably or in tandem; Sennet (2023) points out that “often simple underspecificity will suffice for a charge of ambiguity”.
Egg (2010) gives a detailed categorization of underspecification, grouping instances into four types based on whether the instance’s constituent parts comprise the same semantic value across its readings, and whether it is possible to give a single syntactic analysis for all the readings.
2.2 Semantic Underspecification in NLP
NLP has long studied problems around semantic underspecification. Early work explicitly modeled underspecification by creating formal, symbolic representations of underspecified sentences that captured each sentence’s potential meanings, without generating them (Poesio, 1994; Niehren et al., 1997; Pinkal, 1999, inter alia). NLP systems could then use such representations to make processing sentences more tractable, despite their potentially numerous interpretations Wahlster (2000).
Another line of work focuses instead on identifying or resolving underspecification. Stengel-Eskin et al. (2024), for example, identify the meanings of ambiguous sentences by training a model to map from such sentences to formal representations of their multiple potential meanings. Berzak et al. (2015) train a model to resolve underspecification and determine the correct reading of an ambiguous caption, given an accompanying image. More generally, many classic NLP tasks, such as word sense disambiguation and coreference resolution, involve resolving a word’s meaning in a context where it is underspecified. Where underspecification cannot be resolved, studies have tried to identify or generate clarifications or clarification questions (Roth et al., 2022; Testoni and Fernández, 2024).
Some recent work has addressed ambiguity and underspecification in the context of pre-trained LMs. Considering multi-modal models, Prasad et al. (2023) find that vision-language architectures often struggle with underspecified inputs; specifying inputs improves performance. Pezzelle (2023) reports similarly negative results, discovering that CLIP Radford et al. (2021) sometimes prefers invalid but highly specified captions to valid but underspecified ones. Multi-modal models’ challenges with underspecification concern not only performance but also ethics: Hutchinson et al. (2022) warn that image generation models might rely on social biases to fill in underspecified details.
Related questions have been studied in a uni-modal, text-only context as well, though existing work focuses on ambiguity, rather than underspecification more broadly. For example, Ortega-Martín et al. (2023) and Fantechi et al. (2023) study Chat-GPT’s ability to explicitly identify ambiguity, and report mixed results. Most pertinently, Liu et al. (2023) study how pre-trained LMs process ambiguous sentences, and find that they are unable to use context to infer which potential reading of an ambiguous sentence is correct. In the present work, we aim to expand the existing literature to cover not just ambiguity, but all types of underspecification.
T Phenomenon #S Source
1 Logical Form 35 LAVA
Ellipsis 18 LAVA
2 PP attachment amb. 48 LAVA
VP attachment amb. 60 LAVA
Conjunction amb. 40 LAVA
3 Referential amb. 36 LAVA
Referential amb. 273 WSC
Added compound 774 CLAIRE
Fused head 532 CLAIRE
Implicit reference 216 CLAIRE
Metonymic reference 91 CLAIRE
Table 1: Number of sentences (#S) and source per type (T) and phenomenon of underspecification in DUST. Phenomena containing underspecification without ambiguity are italicized. ‘amb.’ stands for ambiguity.
3 Dataset
To study how LMs deal with semantically underspecified language, and since there currently exists no comprehensive resource on underspecification, we construct DUST, a Dataset of Underspecified Sentences by Type, consisting of 2,123 English underspecified sentences and equally many specified counterparts, based on Egg’s (2010) categorization; see Table 1 for an overview. Below, we discuss the construction of the dataset by type; note that although Egg’s taxonomy includes 4 types of underspecification, we only include 3 in our dataset due to the features of one type (more details below).
Type 1
Egg (2010)’s first type consists of semantically and syntactically homogeneous expressions, i.e., sentences with multiple readings that all share the same syntactic structure, and word/token-level meaning. To cover this type of underspecification, we collect 53 sentences from the Language and Vision Ambiguities (LAVA) dataset (Berzak et al., 2015), a multimodal dataset containing ambiguous sentences and visual data that disambiguated them. For each sentence in this dataset, we created a more specified counterpart using its visual disambiguations. An example of such a sentence is
Andrei approached Danny; Yevgeni, too.
which leaves underspecified whether Yevgeni approached Danny, or was approached by Andrei. A potential specified version of this sentence is
Andrei approached Danny; Andrei approached Yevgeni, too.
Type 2
The second type consists of semantically but not syntactically homogeneous expressions. The multiple readings of such expressions stem from the multiple ways of analyzing the expression’s structure; different analyses can lead to different meanings. The LAVA dataset Berzak et al. (2015) provides 108 sentences containing VP and PP attachment ambiguity and conjunction ambiguity, for example
Andrei looked at the green bag with a telescope.
which leaves ambiguous whether ‘with a telescope’ attaches to ‘the green bag’ or to ‘looked at’. It is thus unclear whether the bag contains a telescope, or was looked at using one. A more specified counterpart of this sentence might be
Andrei looked at the green bag through a telescope.
Type 3
The third type consists of syntactically but not semantically homogeneous expressions, which share a single syntactic structure but do not share the same semantic material in its constituent parts. It is unique in that, besides instances of referential ambiguity, all examples of type 3 are underspecified but not ambiguous. It is thus particularly interesting for our work, as underspecified but not ambiguous expressions are important for LMs to handle, but currently understudied. Examples include deictic expressions and expressions that are underspecified due to missing information.
As examples of this type of underspecification, we first collect 89 sentences from LAVA that contain referential ambiguity or missing information. We then extend our collection with sentences from the CLArifying Insertions from REvision Edits (CLAIRE) dataset Roth et al. (2022), which consists of wikiHow instructional texts, and revisions that clarify the original sentences by inserting additional information. We treat the pre-edit text as underspecified due to missing information, and the post-edit text as more specified. Due to the original authors’ pre-processing, some pre-edit sentences are ungrammatical; we thus score pre-edit sentences’ grammaticality with GRUEN (Zhu and Bhat, 2020) and exclude low-scoring sentences.
We also include the original 273 sentences from the Winograd Schema Challenge (WSC; Levesque et al., 2012) in our dataset as examples of referential ambiguity. For each sentence from this dataset, we crafted a more specified counterpart by changing the gender or plurality of one of the antecedents in the sentence, removing the referential ambiguity. An example sentence of this type is
Don’t spend too much.
which does not specify what should not be spent. A more specified counterpart might be
Don’t spend too much cash.
Type 4
Egg’s fourth and final type, consisting of expressions that are neither syntactically nor semantically homogeneous contains homonymic expressions. These contain a word with multiple meanings; the chosen reading of the homonym can change the meaning of the whole expression. We exclude this type of underspecification from DUST, as it applies mainly to words and phrases. While a word (e.g. “plant”) can have multiple syntactically and semantically distinct readings, most sentences that contain homonyms have readings that rely on the same syntactic structure. For example, “he walked to the bank” contains the homonym “bank”, which could refer to a riverbank of a financial institution, but the syntactic structure of the sentence is the same across both readings. In Appendix A, we report the results and limitations of a preliminary attempt to construct examples falling into this type.
4 Models
We focus on pre-trained autoregressive models, including both older and newer (generally stronger) LMs to consider the influence of general performance improvements on LMs’ underspecification processing. As our experiments require sentence probabilities, we use only openly available models, which provide these. Specifically, we consider the following models, accessed using HuggingFace’s Transformers library Wolf et al. (2020): [^2] See Appendix C for more model details.
  • GPT-2 XL Radford et al. (2019), a 1.5 billion parameter decoder-only LM trained on a 40GB dataset of webpages.
  • FLAN-T5 XXL Chung et al. (2022), a 11 billion parameter encoder-decoder LM. FLAN-T5 is an enhanced version of T5, finetuned in a mixture of language modeling tasks.
  • OPT-13b Zhang et al. (2022), a 13 billion parameter decoder-only transformer model.
  • Llama 2 7b and 13b Touvron et al. (2023), 7 and 13 billion parameter [^3] Due to compute limits, we only use the 13 billion parameter variant of Llama 2 in our second experiment models trained on publicly available online data.
  • Mistral 7b v0.1 Jiang et al. (2023), a 7.3 billion parameter model designed to provide a balance between performance and efficiency.
5 Detecting Semantic Underspecification
We test whether LMs recognize that a sentence is more or less specified by explicitly asking them via a prompt. This task seeks to replicate the process of assessing human speakers’ understanding of this phenomenon by asking them to provide a metalinguistic assessment.
5.1 Experimental Setup
We create inputs of the form “This is an underspecified sentence: [sentence1]. This is its more specified counterpart: [sentence2]”, where [sentence1] and [sentence2] are a pair of sentences from DUST. For each pair, we create a version of the input where the sentences are correctly labeled as under- and more specified and one where their labels are switched.
If the models can recognize underspecification, we would expect that the inputs where the specification labels are correct would receive a higher probability / lower perplexity than the same input with incorrect labels. That is, the input
This is an underspecified sentence: ‘Andrei left the chair with a green telescope’. This is its more specified counterpart: ‘Andrei left the chair on which lay a green telescope’.
should be judged by LMs as more likely than the input where the blue (underspecified) and red (more specified) sentences are switched.
To ensure prompt diversity, and because models may not have been exposed to terminology such as “underspecified” during training, we also use alternate versions of our prompts, where “under-” and “more specified” are replaced by “(un)ambiguous’, or “contain (little/a lot of) (information/detail)”. We also create prompts that reverse the order in which the under- and more specified sentences are presented. [^4] For more example inputs, see Appendix D. In total, we create 33,968 input pairs: 2,123 sentence pairs $\times$ 4 prompt variants $\times$ 4 orders. Then, for each model and input pair, we record whether the model correctly assigns a lower perplexity to the specification-matched input than to its mismatched counterpart.
Figure 2: Proportion of specification-matched inputs that receive lower perplexity than their specification-mismatched counterparts; higher is better. Asterisks (*) indicate performance significantly ($p<0.05$) different from chance (50%). The leftmost column, Overall, reports the proportion computed over the whole dataset.
Sanity check
Our experimental setup aims to determine if models can recognize underspecification when prompted to do so; however, prior work suggests that models may struggle to understand their prompts Webson and Pavlick (2022). So, we first verify the soundness of our setup by using it to test models in an easier domain: sentiment. We gather “very positive” and “very negative” sentences from the SST-5 dataset Socher et al. (2013), and insert them into prompts of the form “This is a positive sentence: [sentence1]. This is a negative sentence: [sentence2]”. All models assign lower perplexity to sentiment-matched inputs with at least 65% and an average of 75% accuracy (see Appendix F for details). This indicates that current LMs can be tested using this experimental setup.
Type Sentence OPT-13B Llama2-7B Mistral 7B
1 Danny left Yevgeni; also Andrei 0.625 0.8125 0.875
Yevgeni approached Andrei; also Danny 0.6875 0.6875 0.8125
2 Andrei looked at Danny moving a blue telescope 0.5 0.625 0.8126
Andrei left the person with a green chair 0.5 0.625 0.6875
3 Pay the fee 0.6875 0.5 0.875
Do this with force, and the can should be smashed 0.4375 0.5 0.5
Table 2: Underspecified sentences and proportion of examples (across prompts and orderings) where the specification-matched input received lower perplexity than its mismatched counterpart; higher is better.
5.2 Results
Newer models perform better
Our results (Figure 2) indicate that some LMs can recognize underspecification. All models besides Flan-T5 do so at a rate significantly higher than chance; Flan-T5’s poor performance may be due to its architecture, i.e., a fine-tuned encoder-decoder model and not a decoder-only model trained on causal language modeling like the others. Stronger models more often prefer specification-matched prompts: Mistral, the best-performing model across the board, achieves an overall accuracy of 0.74. The second-best model, Llama 2, lags behind Mistral by almost 10 accuracy points, with an overall accuracy of 0.65. In turn, this LM outperforms OPT (0.55) and GPT-2 (0.53) by as many accuracy points, indicating a clear ranking between the various models.
Given the models’ high performance in the control experiment with sentiment, the generally lower accuracy observed here is likely due to models’ difficulties in recognizing and identifying underspecification, rather than prompt-related challenges.
Results vary across types
Even for top models, performance across different types of underspecification is not uniform, with gaps of up to 12 accuracy points between the best- and worst-performing types. Mistral, for example, achieves a peak in performance on type 1 (0.85), followed by type 2 (0.79), and type 3 (0.73). For the second best-performing model, Llama 2, type 1 is also the easiest (0.76). However, in contrast with Mistral, types 2 and 3 are equally challenging; Llama 2 achieves similar performance (0.65) on both. These different patterns suggest that the models not only differ in their quantitative ability to perform the task but also in the types of errors they make.
Qualitative analysis
To shed light on the cases where each model succeeds and fails, we conduct a qualitative analysis on a handful of samples, i.e., two underspecified sentences per DUST type, with the best-performing Mistral, Llama 2, and OPT LMs. As reported in Table 2, Mistral is consistently the best-performing model in these examples—it achieves the highest proportion of correctly assigned perplexities. Moreover, the numbers closely align with the patterns reported in Figure 2: for both Mistral and Llama 2, type 1 is the easiest, followed by types 2 and 3. Taking a closer look at the examples, it is worth noting that, (1) for “Pay the fee”, Llama 2 does not recognize it as more underspecified than its counterpart “Pay the licensing fee” (not reported in the table), while OPT and even more Mistral do; (2) for “Do this with force, and the can should be smashed”, all models similarly struggle; none of them recognize the sentence “Do this part with force, and the can should be smashed” (not reported) as its more specified counterpart.
Takeaways and discussion
Overall, our results suggest that modern LMs can moderately identify underspecification if explicitly asked to do so via a prompt. However, our current approach does involve significant prompting, which has several disadvantages when evaluating LMs’ linguistic abilities. Model performance is highly sensitive to the specific prompt used Mizrahi et al. (2024). Moreover, previous work has shown that using meta-linguistic prompts, which explicitly ask for a model’s linguistic judgment, often yields different results than evaluating linguistic tasks using naturalistic data Hu and Levy (2023). While our experiments do not directly ask models if a given sentence is underspecified (we instead compare two versions of the same sentence), our inputs still do not reflect naturalistic data. Our second experiment is motivated by the need to account for this issue.
6 Interpreting Underspecified and Specified Sentences
The results of our first experiment suggest that some LMs may be able to recognize underspecification when asked about it. However, as discussed above, there is no guarantee that the results obtained using metalinguistic prompts are indicative of the actual capabilities of LMs. In the second experiment, we use a more naturalistic setting where underspecification is not mentioned in the prompt.
6.1 Experimental Setup
[…]
Figure 3: Log difference of perplexity of correct and incorrect continuations for underspecified (blue) and more specified (orange) sentences, averaged over prompts. The underspecified condition is not significantly lower than the more specified condition for any model.
We then embed these sentences and continuations in simple templates like “[sentence]. That is, [continuation]”. Given an underspecified sentence, where both continuations are equally plausible, LMs should ideally assign roughly equal perplexity to both. [^6] Underspecified sentences could still have a more frequent “default reading”, making one continuation more likely. For more specified sentences, in contrast, only one interpretation is possible; the corresponding continuation should therefore receive more probability than the incompatible one.
For example, given the underspecified sentence “Danny looked at Andrei with a telescope,” the continuations “That is, Andrei had a telescope,” and “That is, Danny had a telescope,” should be similarly likely. However, given the sentence “Danny looked at Andrei, who had a telescope,” the first continuation would be more likely.
We experiment with connecting the sentence and continuation in different ways; besides inserting “That is,” between them, we also try using no connector (“[sentence]. [continuation]”), and stating a more explicit connection: “[sentence]. Therefore, it is more likely that [continuation1] than [continuation2].”.
We record the absolute value of the difference in the perplexities assigned to each continuation given an underspecified sentence. We expect models will generally have only weak preferences between continuations when given underspecified sentences, leading to smaller absolute differences; specified sentences should have larger differences. For each specified sentence, we also record if LMs prefer the plausible continuation over the implausible.
6.2 Results
Figure 4: Proportion of inputs where the plausible continuation to a specified sentence received lower perplexity; higher is better. Asterisks (*) indicate performance that differs significantly ($p<0.05$) from chance (50%).
LMs do not have stronger preferences toward specified, rather than underspecified sentences
Our results (Figure 3) indicate that the tested models do not interpret underspecified sentences and their specified counterparts correctly: there is no statistically significant difference between the differences in perplexities assigned to continuations of underspecified sentences, and differences in those assigned to continuations of more specified sentences. This suggests that models might incorrectly assign one single interpretation to underspecified sentences, or they might also fail to assign only one interpretation to more specified sentences; a combination of these is also possible.
Sentences and Continuations OPT-13B Llama2-7B Mistral
Andrei looked at Danny holding a yellow bag 20.49 2.26 48.51
Andrei looked at Danny while holding a yellow bag 13.77 2.74 32.61
$\rightarrow$[Andrei / Danny] had a yellow bag
Andrei looked at Danny holding a yellow bag 5.07 2.42 12.81
Andrei looked at Danny while holding a yellow bag 6.03 5.54 13.32
[…]
Andrei looked at Danny holding a yellow bag 1.47 0.68 2.91
Andrei looked at Danny while holding a yellow bag 1.34 0.23 5.84
[…]
Table 3: For each prompt type (base, that is, and more likely), an underspecified sentence and one more specified counterpart, along with the absolute difference in perplexities assigned to the continuations of each. The LM is right when this number is higher for the second, more specified, than the first, underspecified, sentence (bold cases).
The more explicit the prompt, the better the results
Our results (Figure 4) indicate that the type of prompt used heavily influenced the degree to which models preferred the correct, plausible continuation; however, unlike in experiment 1, performance is poor overall. In the base case, where no prompts were used, only Llama 2 7B performs significantly above chance. With the “that is” prompt, model performance improves, though there is no clear trend in which models perform best. It is only with the most extensive prompt that we recover both better performance and the trend from experiment 1, where newer models perform better.
This trend, where more explicit prompts yield better results, is surprising: prior work Hu and Levy (2023) suggested that metalinguistic prompts might more poorly capture LM capabilities, underestimating them. We hypothesize that this could occur because the continuations we crafted are sometimes more like conclusions that follow from the first sentence (as in NLI), and less like genuinely probable continuations. We note, however, that other work has observed that both LMs and humans sometimes default to NLI when given sentence pairs without instructions Webson et al. (2023).
Qualitative analysis
We also examine the degree to which models assign less univocal interpretations to underspecified sentences than their specified counterparts. Table 3 shows the absolute difference in perplexity assigned to the continuations of one under- and one more specified sentence, by prompt and model. On this particular example, the models unanimously assign the more specified sentence a more univocal reading only on the “that is” prompt; on the others, they disagree, echoing the noisiness of our quantitative results.
7 Discussion
Underspecification, much like ambiguity Liu et al. (2023), remains a challenging phenomenon for LMs. Older LMs, such as GPT-2, perform near chance level at recognizing underspecification; newer models, such as Llama 2 and Mistral, perform much better, but still leave ample room for improvement. Processing sentences containing underspecification is an even harder task for LMs. Models seem to fail to recognize when underspecified sentences license continuations that their more specified counterparts do not; moreover, their interpretations of more specified sentences are often incorrect.
The striking difference between the results of our two experiments highlights the importance of carefully choosing a setup when evaluating model capabilities. In the first, metalinguistic prompts elicited good underspecification judgments from high-performing models. But, in surprising contrast to previous work Hu and Levy (2023), testing models’ ability to process underspecification in a more naturalistic setting led to lower performance. This is important: LM use cases involving underspecification will most likely involve processing underspecification, rather than identifying it upon explicit request. Our second experiment may thus be a better indicator of LMs’ practical abilities. However, future work may be needed to compare LM processing of underspecified sentences to results of human studies, which have shown that speakers do have default interpretations of such sentences Kurtzman and MacDonald (1993); Dwivedi (2013).
By introducing DUST and studying underspecification in LMs as distinct from ambiguity, we have taken the first step towards evaluating LMs’ performance on a commonplace but understudied phenomenon that can affect LM behavior. Our findings show that current LMs are limited in their ability to deal with underspecification, especially in genuine communicative scenarios. Hence, a thorough evaluation of the abilities of LMs should include (various facets of) underspecification, unlike current benchmarks, in which ambiguous and underspecified sentences are often systematically excluded. We hope that our research further showcases the relevance of underspecification as a direction of research in the study of language models.
Limitations
DUST is arguably a small dataset, and would benefit from expansion. While we considered existing resources and extracted linguistic data with the desired features from those, future work could expand it by collecting new data via human annotation, generation, or other data-driven approaches. This holds particularly true for type 1, which contains much fewer examples than the other types.
While the present work performs an in-depth evaluation of how LMs behave when faced with semantic underspecification, our research does not explore the inner mechanisms that underlie this capability. We acknowledge that doing so would provide complementary evidence that may be needed to shed full light on the phenomenon.
Our research builds on the formal categorization of semantic underspecification by Egg (2010). While this theoretical framework is both comprehensive and generally suitable for our purposes, we are aware that other theoretical accounts may define the semantic underspecification slightly differently, by including more/less phenomena or by categorizing them according to different criteria. Future work could explore whether and how our findings generalize to other formalizations.
Ethics Statement
While this work presents no serious ethical concerns, a general consideration needs to be made about the use of pre-trained LMs. As is commonly acknowledged, these models should be used with caution as they could perpetuate harmful biases present in their training data. Furthermore, there is a risk that they will generate false or misleading output. In our work, we minimize these risks as we do not use the LMs to generate output, but only to score the plausibility of sentences fed as input. At the same time, we are also aware that some biases may also be present in the linguistic data we used.
Appendix A Type 4 of underspecification
In an attempt to collect sentences containing this type of underspecification, we collect sentences containing homonymic expressions from a sample of English Wikipedia (Mukherjee and Bhattacharyya, 2012). We do so by selecting sentences that contain any homonym from a list of 100 homonyms, selected by Maciejewski and Klepousniotou (2016) based on linguistic principles, dictionary entries, and subjective ratings. We created more specified counterparts by selecting random sentences from the that contain none of the homonyms from the list. We collect a total of 980 sentence pairs; note that unlike other pairs in DUST, these are not minimal pairs. One (partial) sentence of this type is
The elderly fish.
which is underspecified because ‘fish’ can be both a noun or a verb in this context. A more specified counterpart of this (partial) sentence could be
The elderly people fish.
Note, however, that the reading where ‘fish’ is a noun is not a full sentence, and would only be grammatical when placed in a context (e.g. an enumeration) where it is suitable.
Figure 5: Proportion of specification-matched inputs with lower perplexity than their specification-mismatched counterparts; higher is better. Asterisks (*) indicate performance significantly ($p<0.05$) above or below chance (50%).
Running experiment 1 with these sentences included, we get the results shown in Figure 5. We can see that models achieve poor performance on type 4 sentences across models. We hypothesize that this is because that type of underspecification does not consist of minimal pairs; though one item of the pair does include a homonymic expression, this does not guarantee that it is overall less specified than its counterpart. Due to these caveats, we have excluded this type from our dataset.
Appendix B Dataset Information
In this section, we review important licensing and privacy information regarding the datasets composing DUST, as well as DUST itself. DUST is composed of data from 3 datasets.
  • LAVA (Berzak et al., 2015), available at https://web.mit.edu/lavacorpus/, was released with an unclear (potentially open-source) license. LAVA contains images of the authors, which DUST does not include. However, the names used in the sentences in LAVA refer to its authors. These sentences contain no other personally identifiable or offensive content.
  • CLAIRE (Roth et al., 2022), available at https://github.com/acidAnn/claire, is composed of WikiHow articles released under a CC BY-NC-SA 3.0 license, and was itself released under the same license. We did not filter the articles and revisions of which CLAIRE is comprised for personally identifiable or offensive content.
  • The Winograd Schema Challenge dataset (Levesque et al., 2012), available at https://huggingface.co/datasets/winograd_wsc, was released under a CC BY 4.0 license. All sentences were created by experts, and do not contain personally identifiable or offensive content.
DUST is a non-commercial dataset which, like all of its component datasets, is intended for research purposes. We have also provided attribution to the creators of the component datasets, allowing it to be released in accordance with all of the licensing terms of these component datasets. Owing to the WikiHow data contained within, we also release DUST with a CC BY-NC-SA license.
Appendix C Model and Experimental Details
In this section, we provide further information regarding our models and experiments. We use the HuggingFace transformers implementations of all models, available at https://huggingface.co/models. We also use the HuggingFace weights for all models except Llama 2, which must be downloaded separately at https://llama.meta.com/llama-downloads/. To compute perplexities, we use LM-PPL: https://github.com/asahi417/lmppl.
We ran these experiments on compute nodes equipped with Nvidia A100 GPUs (40GB RAM); for all models but Llama 2 13B, one such GPU should suffice. The runtime of our experiments is no more than 5 GPU days.
Appendix D Experiment 1 Prompts
Suppose we have the underspecified sentence ‘Andrei left the chair with a blue telescope’ and the more specified counterpart ‘Andrei left the chair on which lay a blue telescope’. Examples of specification-matched prompts we would then obtain are:
This is an underspecified sentence: ‘Andrei left the chair with a blue telescope’. This is its more specified counterpart: ‘Andrei left the chair on which lay the blue telescope’.
and
This is a sentence that contains a lot of detail: ‘Andrei left the chair on which lay the blue telescope’. This is a sentence that contains little detail: ‘Andrei left the chair with a blue telescope’.
and examples of mismatched prompts are:
This is an ambiguous sentence: ‘Andrei left the chair on which lay the blue telescope’. This is its unambiguous counterpart: ‘Andrei left the chair with a blue telescope’.
and
This is a sentence that contains a lot of information: ‘Andrei left the chair with a blue telescope’. This is its counterpart that contains little information ‘Andrei left the chair on which lay the blue telescope’.
These examples show all variations of phrasing and order of parts
Appendix E Experiment 2 Prompts
Suppose we have the underspecified sentence ‘Andrei looked at Danny moving a green bag’ and the more specified counterpart ‘Andrei looked at Danny who was moving a green bag ’. Examples of specification-matched prompts we would then obtain, from low to high level of prompting, are:
Andrei looked at Danny who was moving a green bag. Danny was moving a green bag.
Andrei looked at Danny who was moving a green bag. Danny was moving a green bag.
Andrei looked at Danny who was moving a green bag. Therefore, it is more likely that Danny was moving a green bag than Andrei was moving a green bag.
and examples of mismatched prompts are
Andrei looked at Danny who was moving a green bag. Andrei was moving a green bag.
Andrei looked at Danny who was moving a green bag. Andrei was moving a green bag.
Andrei looked at Danny who was moving a green bag. Therefore, it is more likely that Andrei was moving a green bag than Danny was moving a green bag.
Appendix F Sentiment perplexity
To test whether the experimental design of experiment 1 functions correctly, we first ran this experiment with sentiment classification instead of recognition of underspecification as a goal. The models were prompted with prompts of the form “prompt1: ’sentence1’. prompt2: ’sentence2’.", where the prompts are of the form "This is a positive/negative sentence" and the sentences are sentences rated ‘very positive’ or ‘very negative’ in the SST-5 dataset (Socher et al., 2013). The results of this can be seen in Figure 6.
Figure 6: Percentage of sentences from experiment 1 with sentiment instead of underspecification where the perplexity of the correct prompts is lower than that of the incorrect prompts, averaged over orders of prompts
Variable Coefficient Standard Error $z$ $P>|z|$ [0.025[0.025[ 0.025 0.975]0.975]0.975 ]
sen. len. -0.155 0.0623 -2.3418 0.2135 -0.2775 -0.0328
avg. AoA. -0.2616 0.2443 -1.046 0.3642 -0.7405 0.2173
avg. conc. 1.7596 0.65 2.6663 0.0262 0.4855 3.034
avg. word freq. -2.9801 2.1483 -1.394 0.1245 -7.1908 1.2307
avg. word len. 0.0367 0.4773 0.062 0.3468 -0.8987 0.972
Table 4: Average regression coefficients averaged over all tested models with sentence length, average age of acquisition Kuperman et al. (2012), average concreteness Brysbaert et al. (2014), average word frequency Norvig (2009) and average word length as independent variables, and whether the perplexity of the correct continuation of the specified counterpart is higher than that of the incorrect continuation as the dependent variable
Appendix G Could models be detecting surface statistics instead of underspecification?
To investigate whether the models’ ability to interpret underspecification as underspecification correlates with some surface-level statistic of the sentences in the dataset, we fit a logistic regression model with surface-level descriptive statistics about each sentence as independent variables and the model ‘correctness’ from Figure 4 as the dependent variable. The results of this can be seen in Table 4.
These results suggest that models are better able to recognize semantic underspecification when the words in the sentence are more concrete. No other surface-level statistic we test shows a significant correlation with the ability of the models to interpret underspecification.
This agrees with intuition: unlike other features like age of acquisition, word frequency or sentence length, concreteness is something humans also associate with (under)specification – for example, when a speaker wants to make things less underspecified, they might say “let’s make things concrete". However, the fact that models do not correctly interpret underspecified sentences when these sentences are abstract in nature does pose a problem, given the fact that certain types of text (e.g. legal texts or product specification documents) can be very abstract while requiring all potential underspecification to be correctly interpreted.